OpenStudy (anonymous):

Chris is diluting 6 gallons of pure apple juice concentrate to a 24% apple juice concentration. How much water should he add? please show all your work.

OpenStudy (anonymous):

A+B=C 6+B=0.24(6)+0.76(B)

OpenStudy (anonymous):

i dunno if that's right, I didn't draw a picture which always helps me

OpenStudy (anonymous):

please can you make a table ?

OpenStudy (anonymous):

wait k let's just work backwards.

OpenStudy (anonymous):

how did you get 0.76

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

total solution is C let's say okay? now since we know that all of the juice mix is going in to the juice we can say that (0.24*C) is 6. tell me if that makes sense to you

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that makes sense? okay then by that you can figure out the total value for C

OpenStudy (anonymous):

or C = 6/.24=25

OpenStudy (anonymous):

can you do the whole problem .step by step .

OpenStudy (anonymous):

Please

OpenStudy (anonymous):

well for material balance problems i usually draw a diagram which helps with everything, but it's not exactly easy on here

OpenStudy (anonymous):

but here your two equations can be this.

OpenStudy (anonymous):

A=(0.24)C B=(0.76)C A+B=C

OpenStudy (anonymous):

3* equations

OpenStudy (anonymous):

now if you can interpret how I say this next part in words it may help you to understand the problem.

OpenStudy (anonymous):

24 percent of the final solution or mixture or whatever you want to call it, will be equal to the total 6 galleons worth of juice concentrate. Originally the conentrate is 100 percent juice concentrate, but by adding water the solution becomes 24 percent juice concentrate so that is all for just the first equation

OpenStudy (anonymous):

basically the 6 galleons of juice concentrate is the 24% value of the total mixture

OpenStudy (anonymous):

so since C = 25 then the amount of water will be 76 percent of the final solution since that's what's remaining. or B = (0.76)*25

OpenStudy (anonymous):

or 19 galleons of water.

OpenStudy (anonymous):

i have a question.how can i get 0.76

OpenStudy (anonymous):

i mean i didn't get that part

OpenStudy (anonymous):

ohhhhhhhh, well it's as simple as this. since 24 percent is juice concentrate, and juice is ONLY made out of juice cocentrate and water, then the water value would be 100%-24% = 76 percent water.

OpenStudy (anonymous):

Thanks .i got it now

OpenStudy (anonymous):

Thank you so much for helping me !

OpenStudy (anonymous):

You are welcome glad I could help!